# Learning Primitive Embodied World Models: Towards Scalable Robotic Learning

### v1

@@ D.1 Dataset Details
Our dataset consists of 7,326 simulated and 11,465 real-world primitives collected using a 5-camera synchronized setup with Deoxys Zhu et al. [2022]. By segmenting long-horizon tasks via keyframe indices, we extracted on average 5.8 primitives per session, achieving up to $29\times$ collection efficiency. As shown in Figure 14, cameras were arranged to maximize coverage of the workspace.
@@ E.1 Base Model Configuration
Our VLM planner is built upon the Qwen2.5-VL-7B-Instruct model Team [2025], a large-scale vision-language model capable of understanding and reasoning over both visual and textual inputs. Qwen-VL integrates a vision encoder based on the Vision Transformer (ViT) architecture to extract image features, which are then mapped into the language model’s embedding space via a cross-modal projector. The underlying LLM, with approximately 7 billion parameters, enables rich semantic understanding and multi-step reasoning, making it well-suited for complex task planning in robotic manipulation.
@@ F.1 Video Generation Model Details
Our base model is DynamiCrafter Xing et al. [2024], which has a 1.4B denoising network.
DynamiCrafter is architected upon the foundation of the open-source Text-to-Video (T2V) model, VideoCrafter Chen et al. [2023, 2024a], and further integrates principles from the Text-to-Image (T2I) model, Stable-Diffusion-v2.1 Rombach et al. [2022a]. This hybrid approach allows DynamiCrafter to leverage the advanced capabilities of both text-conditioned video generation and stable image diffusion, forming a robust framework for animating open-domain images.
@@ G.1 Examples of Training Video Clips
We compare our framework against OpenVLA Kim et al. [2024a], a state-of-the-art end-to-end vision-language-action model that maps raw instructions and observations directly to actions. We evaluate OpenVLA under two settings: (1) Zero-shot: The pre-trained OpenVLA model is deployed directly on our benchmark tasks without any task-specific finetuning. This setting evaluates its raw generalization capability across unseen objects and scenes. (2) Finetune: The model is finetuned on 100 task-specific demonstrations per task using supervised behavior cloning, following the same protocol as in our system’s policy training. This setting represents an upper-bound of OpenVLA’s performance under favorable conditions. We note that OpenVLA performs poorly in the zero-shot setting and fails completely on real-robot deployment, whereas our method maintains strong performance even without task-specific finetuning or retraining of the video model.
@@ H.1 Video Generation Quality
We evaluate visual fidelity using standard metrics on 32-frame action sequences. As shown in Table 6, our method achieves the highest SSIM (0.8126) and PSNR (21.0644), and the lowest TVD (0.0018) and FVD (0.0002), outperforming larger models such as Tesserct (5B) and Hunyuan I2V (13B). This indicates superior structural detail, temporal smoothness, and distributional realism. Despite a slightly higher LPIPS, our outputs exhibit sharper, more dynamic motions that better reflect real robot behavior.
@@ H.2 Physically Accurate Generation
To provide a more comprehensive evaluation of physical fidelity in generated action sequences, we extend our metrics beyond standard video quality indicators (e.g., SSIM, PSNR, LPIPS) to include task- and embodiment-aware measures. In particular, we adopt the evaluation philosophy of PhysicsIQ Motamed et al. [2025] and introduce the first Embodied Physical Consistency Score (EPiCS) tailored for robotic manipulation tasks. EPiCS is a fine-grained, human-in-the-loop metric that assesses whether generated motions respect fundamental physical and kinematic constraints in embodied environments. See a detailed description of EPiCS in Appendix J.
@@ H.3 Supplementary Comparisons
To further contextualize our method’s performance within the broader landscape of vision-based robotic planning, we provide supplementary comparisons on the RLBench benchmark James et al. [2020b] against representative baselines from different paradigms. As shown in Table 7, our approach achieves significantly higher success rates on both Close Box and Open Drawer tasks compared to the VLM-based method VoxPoser Huang et al. [2023] and the action-tokenization approach PerAct Shridhar et al. [2022], where the performance of VoxPoser and PerAct are both extracted from the Colosseum paper Pumacay et al. [2024]. This performance gap highlights the advantage of our video diffusion-based policy generation, which produces temporally coherent and visually grounded action sequences, in contrast to methods that rely on discrete action token prediction or direct VLM-to-action mapping without explicit visual simulation. Notably, our method operates in a zero-shot manner on these tasks—without task-specific fine-tuning of the video generation model—further underscoring its strong generalization capability. These results complement the comprehensive evaluation in Table 1 and reinforce the effectiveness of our framework in handling common, yet non-trivial, manipulation primitives in simulation.

### v2

@@ D.1 Dataset Details
Our dataset consists of 7,326 simulated and 11,465 real-world primitives collected using a 5-camera synchronized setup with Deoxys (Zhu et al., 2022). By segmenting long-horizon tasks via keyframe indices, we extracted on average 5.8 primitives per session, achieving up to $29\times$ collection efficiency. As shown in Figure 14, cameras were arranged to maximize coverage of the workspace.
@@ E.1 Base Model Configuration
Our VLM planner is built upon the Qwen2.5-VL-7B-Instruct model (Team, 2025), a large-scale vision-language model capable of understanding and reasoning over both visual and textual inputs. Qwen-VL integrates a vision encoder based on the Vision Transformer (ViT) architecture to extract image features, which are then mapped into the language model’s embedding space via a cross-modal projector. The underlying LLM, with approximately 7 billion parameters, enables rich semantic understanding and multi-step reasoning, making it well-suited for complex task planning in robotic manipulation.
@@ F.1 Video Generation Model Details
Our base model is DynamiCrafter (Xing et al., 2024), which has a 1.4B denoising network.
DynamiCrafter is architected upon the foundation of the open-source Text-to-Video (T2V) model, VideoCrafter (Chen et al., 2023; 2024a), and further integrates principles from the Text-to-Image (T2I) model, Stable-Diffusion-v2.1 (Rombach et al., 2022a). This hybrid approach allows DynamiCrafter to leverage the advanced capabilities of both text-conditioned video generation and stable image diffusion, forming a robust framework for animating open-domain images.
@@ G.1 Examples of Training Video Clips
We compare our framework against OpenVLA (Kim et al., 2024a), a state-of-the-art end-to-end vision-language-action model that maps raw instructions and observations directly to actions. We evaluate OpenVLA under two settings: (1) Zero-shot: The pre-trained OpenVLA model is deployed directly on our benchmark tasks without any task-specific finetuning. This setting evaluates its raw generalization capability across unseen objects and scenes. (2) Finetune: The model is finetuned on 100 task-specific demonstrations per task using supervised behavior cloning, following the same protocol as in our system’s policy training. This setting represents an upper-bound of OpenVLA’s performance under favorable conditions. We note that OpenVLA performs poorly in the zero-shot setting and fails completely on real-robot deployment, whereas our method maintains strong performance even without task-specific finetuning or retraining of the video model.
@@ H.1 Video Generation Quality
We evaluate visual fidelity using standard metrics on 32-frame action sequences. As shown in Table 6, our method achieves the highest SSIM (0.8126) and PSNR (21.0644), and the lowest TVD (0.0018) and FVD (0.0002), outperforming larger models such as TesserAct (5B) and Hunyuan I2V (13B). This indicates superior structural detail, temporal smoothness, and distributional realism. Despite a slightly higher LPIPS, our outputs exhibit sharper, more dynamic motions that better reflect real robot behavior.
@@ H.2 Physically Accurate Generation
To provide a more comprehensive evaluation of physical fidelity in generated action sequences, we extend our metrics beyond standard video quality indicators (e.g., SSIM, PSNR, LPIPS) to include task- and embodiment-aware measures. In particular, we adopt the evaluation philosophy of PhysicsIQ (Motamed et al., 2025) and introduce the first Embodied Physical Consistency Score (EPiCS) tailored for robotic manipulation tasks. EPiCS is a fine-grained, human-in-the-loop metric that assesses whether generated motions respect fundamental physical and kinematic constraints in embodied environments. See a detailed description of EPiCS in Appendix J.
@@ H.3 Supplementary Comparisons
To further contextualize our method’s performance within the broader landscape of vision-based robotic planning, we provide supplementary comparisons on the RLBench benchmark (James et al., 2020b) against representative baselines from different paradigms. As shown in Table 7, our approach achieves significantly higher success rates on both Close Box and Open Drawer tasks compared to the VLM-based method VoxPoser (Huang et al., 2023) and the action-tokenization approach PerAct (Shridhar et al., 2022), where the performance of VoxPoser and PerAct are both extracted from the Colosseum paper (Pumacay et al., 2024). This performance gap highlights the advantage of our video diffusion-based policy generation, which produces temporally coherent and visually grounded action sequences, in contrast to methods that rely on discrete action token prediction or direct VLM-to-action mapping without explicit visual simulation. Notably, our method operates in a zero-shot manner on these tasks—without task-specific fine-tuning of the video generation model—further underscoring its strong generalization capability. These results complement the comprehensive evaluation in Table 1 and reinforce the effectiveness of our framework in handling common, yet non-trivial, manipulation primitives in simulation.